Two loudspeakers in a 20 ?C room emit 660 Hz sound waves along the x -axis. Part A What is the smallest distance between the speakers for which the interference of the sound waves is destructive?
Answer #1

given frequency f = 660 Hz

speed of sound v = 343 m/s

from the relation

lambda*f = v

lambda = v/f = 343/660 = 0.519 m

for distructive interference path differenc eis = d = lambda/2

d = 0.519/2 = 0.259 m